# Subjective Behaviors and Preferences in LLM: Language of Browsing

### v1

@@ Abstract
Sai Sundaresan  Harshita Chopra [^2] Work done while authors were at Adobe Research  Atanu R. Sinha Koustava Goswami  Nagasai Saketh Naidu  Raghav Karan [^2] Work done while authors were at Adobe Research  N Anushka [^2] Work done while authors were at Adobe Research Adobe Research atr@adobe.com
@@ 1 Introduction
Users’ everyday interactions with websites and apps occur in an idiosyncratic manner, generating a language of sequence of webpages (or, pages), termed language of browsing. Fig 1 shows an example of this. This language exemplifies subjective behaviors and preferences, which are heterogeneous and crucially do not conform to oracle like behavior nor oracle-preference. Even if a single LM produces good average performance in generation across users, the variance in performance across users may be weak, making user level alignment weak. Thus, alignment with some ideal oracle or with the average across users, is less meaningful. The language of browsing is ubiquitous in the form of pageurl-logs that online firms collect. We study users’ heterogeneity of behaviors to examine the effectiveness of a single LM. Domain adaptation of LM addresses aspects of heterogeneity in domains and tasks, but none of these works tackle the important problem of heterogeneity in users’ subjective behaviors and preferences, embodied in the sequential pages browsed. We learn the language intrinsic to such sequences, using decoder LM. For online firms, prediction of outcomes such as user conversion is of utmost importance. We thus consider three goals: (a) average performance in generation of pages, (b) variance of performance in generation of pages across users, and (c) performance on outcome prediction, where (b) and (c) are unattended in LM prior art. We find that small LMs, pretrained using a page-level tokenizer, outperform GPT4o with prompting, and fine-tuned versions of LLaMa-3-8B, Mistral-7B, Gemma-7B. The small LMs, are then finetuned clusterwise through our proposed HeTLM (Heterogeneity aware Training of Language Model), where both the number of clusters $K$ and assignment of users to clusters are endogenously determined as finetuning progresses. The average performance in generation across users for HeTLM is better than the single LM of larger size in the same family, and the variance in performance is lower. Also, HeTLM performs well on prediction of outcomes available in the data–add to cart and purchase. These show that HeTLM improves performance for both page generation and outcome prediction. Our contributions are: (i) Drawing attention to Language of Browsing for LLMs; (ii) Showing that a small model can outperform a large model for outcome prediction with effective pretraining; (iii) Introducing heterogeneity of users’ subjective behaviors and preferences in LLMs; (iv) Offering a network architecture to train LLM clusterwise.
@@ 5.1 Evaluation Metrics for Step 1
All evaluations are performed by comparing page generation in next session against the pages in the Actual next session (e.g., Fig. 1). In Step 1, our first level comparison, we use 6 evaluation metrics. In Step 2, we use 14 additional metrics to deep dive for detailed comparisons. Two overarching evaluation objectives relate to Page Generation, and Outcome Prediction. To address our thesis of clusterwise LMs, we introduce Variance in Page Generation metric across users, as an additional evaluation metric, described in 5.2.2.
@@ 5.1.1 Page Generation - Mean across Users
[IoA] Intersection over Actual: Ratio of number of correctly generated pages to that of actual pages; same as, $1$ minus False Negative.
[IoP] Intersection over Predicted (Generated): Ratio of number of correctly generated pages to that of generated pages; same as, $1$ minus False Positive.
@@ 6 Experiment Results: Dataset I
6.1 Results: Step 1
Small LMs, OPT-350M, QWEN-2.5-500M and SmolLM2-360M, are pretrained on Dataset I, while large LMs, Llama3-8B, Mistral-7B and Gemma-7B, are finetuned on the same data. Per Table 1, small LMs outperform large LMs, namely, GPT-4o-200B+, LLaMa2-7B, Llama3-8B, Mistral-7B, Gemma-7B. Step 2 HeTLM clusterwise experiments are done with small LMs.
6.2 Results: Step 2
Model
Outcome
Prediction
Page Gen
Mean
Page Gen
Var
Overall
Composite
Compared with OPT-2.7B
OPT-350M - Combined
Kmeans, K=6 0.0 0.333 0.75 0.25
HeTLM ($\alpha$=2, $\beta$=1) 0.3 0.167 0.75 0.35
HeTLM ($\alpha$=5, $\beta$=9) 0.5 0.5 1.0 0.6
Compared with QWEN-2.5-7B
QWEN-2.5-500M - Combined
HeTLM ($\alpha$=5, $\beta$=9) 1.0 0.833 0.25 0.8
6.3 BLEU and BERT metrics
@@
Model NN
HR-var
IoA-var
IoP-var
IoU-var
OPT 2.7B 5253 0.152 0.121 0.119 0.109
OPT-350M: Kmeans, K=6
Cluster 1 438 0.136 0.088 0.076 0.053
Cluster 2 1359 0.209 0.047 0.031 0.015
Cluster 3 638 0.158 0.068 0.06 0.031
Cluster 4 330 0.19 0.155 0.148 0.129
Cluster 5 573 0.211 0.087 0.108 0.075
Cluster 6 1915 0.163 0.069 0.07 0.044
Combined 5253 0.182 0.077 0.074 0.048
OPT-350M: HeTLM ($\alpha$=2, $\beta$=1)
Cluster 1 948 0.163 0.094 0.09 0.071
Cluster 2 319 0.105 0.138 0.122 0.129
Cluster 3 2875 0.16 0.1 0.102 0.083
Cluster 4 277 0.215 0.102 0.133 0.091
Cluster 5 759 0.17 0.061 0.072 0.037
Combined 5253 0.162 0.101 0.103 0.082
OPT-350M: HeTLM ($\alpha$=5, $\beta$=9)
Cluster 1 305 0.107 0.137 0.129 0.144
Cluster 2 4797 0.148 0.096 0.097 0.082
Cluster 3 151 0.245 0.143 0.185 0.137
Combined 5253 0.15 0.103 0.107 0.092
QWEN-2.5-7B 5253 0.211 0.064 0.067 0.036
QWEN-2.5-500M: HeTLM ($\alpha$=5, $\beta$=9)
Cluster 1 3101 0.154 0.125 0.103 0.098
Cluster 2 316 0.113 0.143 0.135 0.151
Cluster 3 1673 0.161 0.123 0.115 0.109
Cluster 4 163 0.241 0.14 0.170 0.128
Combined 5253 0.158 0.129 0.117 0.111
@@ A.1 Theoretical Basis of HeTLM
Following Lee and van der Schaar (2020), we formulate the problem of identifying user clusters with similar browsing behaviors and preferences as a predictive clustering problem. Let $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$ be random variables for input browsing sessions and output next session pages with a joint distribution $p_{XY}$, where $\mathcal{X}$ is the browsing session space and $\mathcal{Y}$ is the next session page space.
@@
We define a cluster as $\mathcal{C}(k)=\{x^{n}|t\,~s^{n}=k\}$ for $k\in\{1,...,K\}$ where $s^{n}\in\{1,...,K\}$ is the cluster assignment for user $n$. This allows us to flexibly update the cluster assignment to which a user belongs as the cluster representations are updated over time.
Let $S$ be a random variable for the cluster assignment. We want to find an optimal partitioning of users into $K$ clusters such that the difference in the next session distribution conditioned on the input sessions $X$ and conditioned on the cluster assignment $S$ is minimized, while optimizing the number of clusters $K$. This can be achieved by minimizing the following Kullback-Leibler (KL) divergence:
@@
The optimization problem is non-trivial, and we estimate this objective function through the LLM loss ($\mathcal{L}_{1}$) described in Sec. 4.5. Minimizing $\mathcal{L}_{1}$ is equivalent to minimizing the KL divergence in Eq. 5 since the former term of the KL divergence is independent of our optimization procedure. To avoid trivial solutions in this unsupervised setting—such as all embeddings collapsing to a single point or the selector assigning equal probability to all clusters regardless of the input sequence—we introduce two auxiliary loss functions, $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$, as detailed in Sec. 4.5.
A.2 Dataset I: Additional Experiments
Extensive results for Dataset I are presented in the main paper. For completeness, in Appendix, we present results for HeTLM for another small LM, namely, SmolLM2-360M.
A.2.1 Experimental Results: HeTLM for SmolLM2-360M
For small LM, SmolLM2-360M, we present results for clusterwise training. Note that we did not have a large variant of this LM and hence could not make a comparison of the HeTLM with the large variant. However, the clusterwise HeTLM results here show considerable variations across clusters in all metrics of Tables 7, 8, and 9. This calls attention to the importance of clusterwise training to achieve better alignment of users’ behaviors.
@@ A.2.1 Experimental Results: HeTLM for SmolLM2-360M
A.3 Experiments: Dataset II
A.3.1 Dataset II
This dataset is public as well Kechinov (2019). Train : Test split = 1M : 10,000 samples. Number of unique pages = 17,310. The dataset has millions of consumers; after applying filters for excluding very short sessions and incomplete data, we randomly select 1.01 million samples, which are split into train and test sets. Outcomes available in the pageurls are: Cart (Add to Cart, in Dataset II) and Purchase, these Outcome pages are generated by the LM and also used as target labels to evaluate Outcome prediction against ground truth of Actual (next) session.
A.3.2 Experimental Results: Dataset II
We present comparison of HeTLM with OPT-2.7B In Page Generation, comparing the respective Combined row in Table 11 with OPT-2.7B shows that except Val-P, in all other metrics Kmeans and HeTLM perform slightly better. For HeTLM, the variations in number of clusters and size of clusters and their differences in metrics justify endogenous clustering based training. In Outcome Prediction (Table 10) across its 10 metrics HeTLM has an edge over OPT-2.7B. For Variance in Page Generation (Table 12) both models perform similarly. Reviewing Overall Composite metric in Table 13 finds HeTLM outperforming OPT-2.7B, as the values [0.7, 0.6] are >> 0.5. APPENDIX contains additional results from Dataset II for more variations of HeTLM and Kmeans.
@@ A.3.2 Experimental Results: Dataset II
Table 12: Dataset II. Variance in Page Generation across users.
@@ A.4 Experiment Details
A.4.1 Small Model Pretraining
Hyperparameter Value
Base Model
OPT-350M
QWEN-2.5-500M
SmolLM2-360M
Tokenizer Type Custom (page level)
Learning Rate 1e-5
Weight Decay 0.01
Batch Size 32
Maximum Sequence Length 512
Validation Set Size 0.05
Number of Epochs 100
Gradient Accumulation Steps 4
Warmup Ratio 0.05
Table 14: Small LM Pretraining Hyperparameters
A.4.2 HeTLM Training
Table 15 presents the hyperparameters used for training our proposed Heterogeneity-aware Training of Language Models (HeTLM). These parameters control the Actor-Critic architecture and the weight of various loss components.
@@ A.4.2 HeTLM Training
A.4.3 Large Model LoRA Finetuning
@@
A.4.4 Inference Time Comparison
Table 17 compares total inference time for different model configurations on Dataset I (10k samples). All measurements were conducted under identical hardware settings with maximum GPU utilization.
@@ A.4.4 Inference Time Comparison
Notably, our HeTLM approach achieves faster inference times than larger single models while having better performance (Sec. 6).

### v2

@@ Abstract
Sai Sundaresan1  Harshita Chopra2222Work done while authors were at Adobe Research  Atanu R. Sinha1 Koustava Goswami1  Nagasai Saketh Naidu1  Raghav Karan1222Work done while authors were at Adobe Research  N Anushka1222Work done while authors were at Adobe Research 1Adobe Research 2University of Washington, Seattle atr@adobe.com
@@ 1 Introduction
Users’ everyday interactions with websites and apps occur in an idiosyncratic manner, generating a language of sequence of webpages (or, pages), termed language of browsing. Fig 1 shows an example of this. This language exemplifies subjective behaviors and preferences, which are heterogeneous and crucially do not conform to oracle like behavior nor oracle-preference. Even if a single LM produces good average performance in generation across users, the variance in performance across users may be weak, making user level alignment weak. Thus, alignment with some ideal oracle or with the average across users, is less meaningful. The language of browsing is ubiquitous in the form of pageurl-logs that online firms collect. We study users’ heterogeneity of behaviors to examine the effectiveness of a single LM. Domain adaptation of LM addresses aspects of heterogeneity in domains and tasks, but none of these works tackle the important problem of heterogeneity in users’ subjective behaviors and preferences, embodied in the sequential pages browsed. We learn the language intrinsic to such sequences, using decoder LM. For online firms, prediction of outcomes such as user conversion is of utmost importance. We thus consider three goals: (a) average performance in generation of pages, (b) variance of performance in generation of pages across users, and (c) performance on outcome prediction, where (b) and (c) are unattended in LM prior art.
We find that small LMs, pretrained using a page-level tokenizer, outperform GPT4o with prompting, and fine-tuned versions of LLaMa-3-8B, Mistral-7B, Gemma-7B. The small LMs, are then finetuned clusterwise through our proposed HeTLM (Heterogeneity aware Training of Language Model), where both the number of clusters $K$ and assignment of users to clusters are endogenously determined as finetuning progresses. The average performance in generation across users for HeTLM is better than the single LM of larger size in the same family, and the variance in performance is lower. Also, HeTLM performs well on prediction of outcomes available in the data–add to cart and purchase. These show that HeTLM improves performance for both page generation and outcome prediction.
Our contributions are: (i) Drawing attention to Language of Browsing for LLMs; (ii) Showing that a small model can outperform a large model for outcome prediction with effective pretraining; (iii) Introducing heterogeneity of users’ subjective behaviors and preferences in LLMs; (iv) Offering a network architecture to train LLM clusterwise.
@@ 5.1 Evaluation Metrics for Step 1
All evaluations are performed by comparing page generation in next session against the pages in the Actual next session (e.g., Fig. 1). We use a broad set of metrics to support three key aspects of our approach. (1) Variance metrics in page generation are introduced to capture variability across users—an aspect overlooked in prior work but central to our premise of subjectivity in user heterogeneity. This aligns with our thesis of clusterwise LMs and is described in 5.2.2. (2) Page generation metrics are aligned with prior work on language models and allow standard benchmarking. (3) Outcome prediction metrics (Accuracy, Precision, Recall, F1) are widely used in prior art and are essential for evaluating business impact. We further disaggregate these by outcome type (e.g., cart vs. purchase) to give the reader a more complete view of model behavior. Rather than reducing to a minimal set, we provide comprehensive metrics so readers can interpret results according to their own priorities. To aid in comparison across models, we also include composite metrics in Sec. 5.3. In Step 1, our first level comparison, we use 6 evaluation metrics. In Step 2, we use 14 additional metrics to deep dive for detailed comparisons.
@@ 5.1.1 Page Generation - Mean across Users
[IoA] Intersection over Actual: Ratio of number of correctly generated pages to that of actual pages; same as, 11 minus False Negative.
[IoP] Intersection over Predicted (Generated): Ratio of number of correctly generated pages to that of generated pages; same as, 11 minus False Positive.
@@ 6 Experiment Results: Dataset I
Model NN
HR-var
IoA-var
IoP-var
IoU-var
OPT 2.7B 5253 0.152 0.121 0.119 0.109
OPT-350M: Kmeans, K=6
Cluster 1 438 0.136 0.088 0.076 0.053
Cluster 2 1359 0.209 0.047 0.031 0.015
Cluster 3 638 0.158 0.068 0.06 0.031
Cluster 4 330 0.19 0.155 0.148 0.129
Cluster 5 573 0.211 0.087 0.108 0.075
Cluster 6 1915 0.163 0.069 0.07 0.044
Combined 5253 0.182 0.077 0.074 0.048
OPT-350M: Kmeans, K=2
Combined 5253 0.174 0.096 0.092 0.071
OPT-350M: HeTLM ($\alpha$=2, $\beta$=1)
Cluster 1 948 0.163 0.094 0.09 0.071
Cluster 2 319 0.105 0.138 0.122 0.129
Cluster 3 2875 0.16 0.1 0.102 0.083
Cluster 4 277 0.215 0.102 0.133 0.091
Cluster 5 759 0.17 0.061 0.072 0.037
Combined 5253 0.162 0.101 0.103 0.082
OPT-350M: HeTLM ($\alpha$=5, $\beta$=9)
Cluster 1 305 0.107 0.137 0.129 0.144
Cluster 2 4797 0.148 0.096 0.097 0.082
Cluster 3 151 0.245 0.143 0.185 0.137
Combined 5253 0.15 0.103 0.107 0.092
QWEN-2.5-7B 5253 0.211 0.064 0.067 0.036
QWEN-2.5-500M: HeTLM ($\alpha$=5, $\beta$=9)
Cluster 1 3101 0.154 0.125 0.103 0.098
Cluster 2 316 0.113 0.143 0.135 0.151
Cluster 3 1673 0.161 0.123 0.115 0.109
Cluster 4 163 0.241 0.14 0.170 0.128
Combined 5253 0.158 0.129 0.117 0.111
@@
6.1 Results: Step 1
Small LMs, OPT-350M, QWEN-2.5-500M and SmolLM2-360M, are pretrained on Dataset I, while large LMs, Llama3-8B, Mistral-7B and Gemma-7B, are finetuned on the same data. Two variations of Llama3-8B are used - LORA finetuned and full finetuned. Per Table 1, small LMs outperform large LMs, namely, GPT-4o-200B+, LLaMa2-7B, Llama3-8B, Mistral-7B, Gemma-7B. Step 2 HeTLM clusterwise experiments are done with small LMs.
6.2 Results: Step 2
In addition to results shown for Kmeans, K=6, we also show results for K=2 (decided based on Silhouette Coefficient). See Tables 25. We find that K=2 produces worse performance than K=6 for Kmeans, where choice of number of clusters is fixed exogenously. However, notably for HeTLM, choice of number of clusters is endogenous and HeTLM outperforms Kmeans.
Model
Outcome
Prediction
Page Gen
Mean
Page Gen
Var
Overall
Composite
Compared with OPT-2.7B
OPT-350M - Combined
Kmeans, K=6 0.0 0.333 0.75 0.25
Kmeans, K=2 0.1 0.0 0.5 0.15
HeTLM ($\alpha$=2, $\beta$=1) 0.3 0.167 0.75 0.35
HeTLM ($\alpha$=5, $\beta$=9) 0.5 0.5 1.0 0.6
Compared with QWEN-2.5-7B
QWEN-2.5-500M - Combined
HeTLM ($\alpha$=5, $\beta$=9) 1.0 0.833 0.25 0.8
@@ 6.2 Results: Step 2
6.3 BLEU and BERT metrics
Model
Outcome
Prediction
Page Gen
Mean
Page Gen
Var
Overall
Composite
Compared with OPT-2.7B
OPT-350M - Combined
HeTLM ($\alpha$=1, $\beta$=0) 0.1 0.0 0.75 0.2
HeTLM ($\alpha$=1, $\beta$=3) 0.4 0.167 0.75 0.4
HeTLM ($\alpha$=9, $\beta$=0) 0.1 0.167 0.75 0.25
HeTLM ($\alpha$=9, $\beta$=3) 0.1 0.0 0.75 0.2
Table 7: Dataset I. Ablation. Composite metrics of Sec. 5.3. Ablation with OPT-350M, Pre-training, by varying $\alpha$, $\beta$.
6.4 Ablation: OPT-350M
We ran two ablations to examine the effect of the sharpness of the selector probability distribution ($\alpha$) and the presence of cluster separation ($\beta$). The hyperparameters $\alpha$, $\beta$ are selected empirically and these are dataset and LM specific. We provide a snapshot of composite metrics in Table 7. By comparing with results of HeTLM for OPT-350M in Table 5, we find that the proposed architecture with non-zero $\alpha$=5, $\beta$=9 performs appreciably better than suppressing either $\alpha$, or $\beta$. General trends may be identified with a more indepth analysis which is beyond the scope of this paper.
@@ A.1 Theoretical Basis of HeTLM
Following Lee and van der Schaar (2020), we formulate the problem of identifying user clusters with similar browsing behaviors and preferences as a predictive clustering problem. Let $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$ be random variables for input browsing sessions and output next session pages with a joint distribution $p_{XY}$, where $\mathcal{X}$ is the input session space and $\mathcal{Y}$ is the next session page space.
@@
We define a cluster as $\mathcal{C}(k)=\{x^{n}|t,~s^{n}=k\}$ for $k\in\{1,...,K\}$ where $s^{n}\in\{1,...,K\}$ is the cluster assignment for user $n$. This allows us to flexibly update the cluster assignment to which a user belongs as the cluster representations are updated over time.
Let $S$ be a random variable for the cluster assignment. We want to find an optimal partitioning of users into $K$ clusters such that the difference in the next session distribution conditioned on the input sessions $X$ and the cluster assignment $S$ is minimized, while optimizing the number of clusters $K$. This can be achieved by minimizing the following Kullback-Leibler (KL) divergence:
@@
The optimization problem is non-trivial, and we estimate this objective function through the LLM loss ($\mathcal{L}_{1}$) described in Sec. 4.5. Minimizing $\mathcal{L}_{1}$ is equivalent to minimizing the KL divergence in Eq. 5 since the former term is independent of our optimization procedure. To avoid trivial solutions in this unsupervised setting, such as all embeddings collapsing to a single point or the selector assigning equal probability to all clusters regardless of the input sequence, we introduce two auxiliary loss functions, $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$, as detailed in Sec. 4.5.
A.2 Experimental Results: HeTLM for SmolLM2-360M
For small LM, SmolLM2-360M, we present results for clusterwise training. The clusterwise HeTLM results here show considerable variations across clusters in all metrics of Tables 8, 9, and 10. This calls attention to the importance of clusterwise training to achieve better alignment of users’ behaviors.
@@ A.2 Experimental Results: HeTLM for SmolLM2-360M
A.3 Dataset II: Experiments
This dataset is public as well Kechinov (2019). Train : Test split = 1M : 10,000 samples. Number of unique pages = 17,310. The dataset has millions of consumers; after applying filters for excluding very short sessions and incomplete data, we randomly select 1.01M samples. Outcomes available in the pageurls are: Cart (Add to Cart, in Dataset II) and Purchase, these Outcome pages are generated by the LM and also used as target labels to evaluate Outcome prediction.
A.3.1 Experimental Results: Dataset II
We present comparison of HeTLM with OPT-2.7B. In Page Generation, comparing the respective Combined row in Table 12 with OPT-2.7B shows that except Val-P, in all other metrics Kmeans and HeTLM perform slightly better. For HeTLM, the variations in number of clusters and size of clusters and their differences in metrics justify endogenous clustering based training. In Outcome Prediction (Table 11) across its 10 metrics HeTLM has an edge over OPT-2.7B. For Variance in Page Generation (Table 13) both models perform similarly. Reviewing Overall Composite metric in Table 14 finds HeTLM outperforming OPT-2.7B, as the values [0.7, 0.6] are >> 0.5. APPENDIX contains additional results from Dataset II for more variations of HeTLM and Kmeans.
@@ A.3.1 Experimental Results: Dataset II
Table 13: Dataset II. Variance in Page Generation across users lower is better.
@@ A.4 Experimental Setup Details
A.4.1 Prompt Templates
The zero-shot setup used a basic instruction prompt: "A user’s website browsing sequence of pages over multiple sessions is given as Input. INPUT SESSIONS [BOS] denotes beginning of session, [EOS] denotes end of session. Based on activity in the given input sessions, predict the next session as a sequence of pages highly likely to be visited by the user". For the few-shot setup we add 3 examples of input-output pairs based on cosine similarity of SBERT embeddings with the user’s input sessions: "Learn from similar users browsing sequences given below: INPUT-OUTPUT PAIRS".
A.4.2 Small Model Pretraining
Hyperparameter Value
Base Model
OPT-350M
QWEN-2.5-500M
SmolLM2-360M
Tokenizer Type Custom (page level)
Learning Rate 1e-5
Weight Decay 0.01
Batch Size 32
Maximum Sequence Length 512
Validation Set Size 0.05
Number of Epochs 100
Gradient Accumulation Steps 4
Warmup Ratio 0.05
Table 16: Small LM Pretraining Hyperparameters
A.4.3 HeTLM Training
Table 15 details the hyperparameters used for training our proposed Heterogeneity-aware Training of Language Models (HeTLM). These parameters control the Actor-Critic architecture and the weight of various loss components.
A.4.4 Large Model LoRA Finetuning
@@
A.4.5 Inference Time Comparison
Table 18 compares total inference time for different model configurations on Dataset I (10k samples). All measurements were conducted under identical hardware settings with maximum GPU utilization. Notably, our HeTLM approach achieves faster inference times than larger single models while having better performance (Sec. 6).